# MCGI: Manifold-Consistent Graph Indexing for Billion-Scale Disk-Resident Vector Search

### v1

@@
Abstract.
[…]
Approximate Nearest Neighbor Search, Disk-resident Indexing, Local Intrinsic Dimensionality
copyright: noneccs: Information systems Top-k retrieval in databases
1. Introduction
@@
The advent of Large Language Models (LLMs) (brown2020language; touvron2023llama) has fundamentally transformed the landscape of information retrieval and knowledge management. To address the inherent limitations of LLMs, such as hallucinations (ji2023survey) and knowledge cutoff dates, Retrieval-Augmented Generation (RAG) (lewis2020retrieval) has emerged as a critical architectural paradigm. RAG relies heavily on the ability to retrieve semantically relevant context from massive corpora in real-time, typically via dense vector representations (karpukhin2020dense). This dependency has placed Approximate Nearest Neighbor Search (ANNS) at the core of modern data infrastructure, demanding vector indices that can scale to billion-point datasets (simhadri2022neurips) while maintaining low latency and high recall under strict production constraints.
@@
State-of-the-art ANNS solutions have largely converged on graph-based indices, with DiskANN (Vamana) (subramanya2019diskann) being a representative example for SSD-resident workloads. These algorithms typically employ greedy routing on a proximity graph to navigate from an entry point to the query target. While such methods exhibit exceptional performance on standard benchmarks like SIFT1M (jegou2011product) (128 dimensions), their efficiency degrades significantly in high-dimensional spaces, such as GIST1M (960 dimensions). This degradation is often attributed to the curse of dimensionality (bellman1957dynamic), where the distance contrast diminishes, and the Euclidean shortest path on the graph diverges from the geodesic path on the underlying data manifold. We refer to this phenomenon as the Euclidean-Geodesic mismatch. When the routing algorithm ignores the intrinsic geometry of the data, it performs excessive backtracking and disk I/O, rendering the search inefficient, particularly problematic in high-recall regimes where production systems usually expect Recall@10 $\geq$ 95%.
@@
Our key insight is that high-dimensional real-world data is rarely uniformly distributed. Instead, it typically adheres to the Manifold Hypothesis (tenenbaum2000global; roweis2000nonlinear), residing on lower-dimensional structures embedded within the ambient space. Consequently, the search difficulty is not uniform across the dataset but is modulated by the Local Intrinsic Dimensionality (LID) (NIPS2004_74934548). In regions where the data manifold is flat (low LID), greedy routing is effective; however, in regions with high curvature or complex topology (high LID), standard greedy strategies fail to identify the correct descent direction. We argue that an optimal indexing strategy must be manifold-aware, dynamically allocating computational resources based on the local geometric complexity. To address these challenges, we introduce Manifold-Consistent Graph Indexing (MCGI), a geometry-aware disk-based indexing architecture designed to align Euclidean search with the underlying data manifold. By integrating LID estimation directly into the routing logic, MCGI adapts its traversal strategy to the local topology of the data.
@@
  • We establish a theoretical method linking local intrinsic dimensionality to graph navigability, offering a justification for adaptive beam search on non-Euclidean manifolds.
@@
  • We develop a lightweight adaptive routing algorithm that dynamically modulates the search budget based on real-time geometric analysis. This design eliminates the dependency on static, manually tuned hyperparameters that limits the adaptability of existing methods.
[…]
2. Related Work
Vector Indexing Paradigms.
While traditional sparse retrieval methods like BM25 (robertson2009probabilistic) rely on lexical matching, the surge of neural networks has shifted the focus to dense vector retrieval. In the memory-resident regime, graph-based indices, particularly Hierarchical Navigable Small World (HNSW) (malkov2018efficient), have established state-of-the-art performance by enabling logarithmic complexity scaling. However, the high memory consumption of HNSW poses challenges for billion-scale datasets. To mitigate this, disk-based approaches have emerged. DiskANN (Vamana) (subramanya2019diskann) adapts the graph topology for SSDs by relaxing sparsity constraints to maximize neighborhood coverage. Concurrently, methods like SPANN (chen2021spann) argue against pure graph traversal on disk due to random I/O latency, advocating instead for an inverted index (IVF) structure combined with centroid-based routing. Despite their differences, both DiskANN and SPANN, as well as their predecessors like NSG (fu2019nsg), are predominantly evaluated on standard benchmarks such as SIFT and DEEP with moderate dimensionality (96 to 128). They largely rely on static routing parameters or centroid layouts that do not explicitly account for the local intrinsic dimensionality. In contrast, MCGI distinguishes itself by abandoning static routing configurations in favor of a geometry-aware strategy. By dynamically modulating the search budget based on estimated LID, our method aligns the graph traversal with the underlying manifold structure, thereby overcoming the efficiency bottlenecks that hinder rigid indexing schemes in high-dimensional spaces.
High-Dimensional Indexing.
Early approaches relied on space-partitioning trees. The KD-tree (bentley1975multidimensional) divides the space using axis-aligned hyperplanes, while the R-tree (guttman1984r) utilizes hierarchical bounding rectangles. However, these strict partitioning schemes suffer from the curse of dimensionality, typically degrading to linear scan performance when the dimension exceeds 20. To address this, approximate methods like Locality-Sensitive Hashing (LSH) (indyk1998approximate; datar2004locality) were introduced, offering sub-linear search time guarantees. Yet, achieving high recall with LSH often requires maintaining significant redundancy via multiple hash tables, resulting in excessive storage overhead. Another line of work involves subspace quantization, exemplified by Product Quantization (PQ) (jegou2011product) and Optimized PQ (OPQ) (ge2013optimized), which decompose high-dimensional vectors into lower-dimensional subspaces for compression. Additionally, randomized structures like RP-Trees (dasgupta2008random) attempt to adapt to the data geometry via random projections. MCGI differs from these approaches by retaining the connectivity benefits of graph traversal while avoiding the aggressive quantization loss or the storage redundancy of hashing methods.
Intrinsic Dimensionality.
Theoretical analysis of nearest neighbor search often relies on characterizing the intrinsic difficulty of the dataset. Fundamental concepts such as the doubling dimension (gupta2003bounded) and expansion dimension (karger2002finding) provide asymptotic bounds on search complexity in growth-restricted metrics. Bridging theory and practice, Levina and Bickel (NIPS2004_74934548) introduced maximum likelihood estimators for Local Intrinsic Dimensionality (LID), enabling robust estimation on real-world data. While subsequent works have utilized LID for tasks such as query hardness prediction (he2012difficulty) or detecting adversarial examples (ma2018characterizing), these applications are typically passive, utilizing LID primarily for post-hoc analysis or pre-query estimation without altering the underlying index structure. MCGI diverges from this paradigm by employing LID as an active control signal. By dynamically modulating the graph traversal parameters based on local geometry, our method transforms LID from a descriptive metric into a prescriptive mechanism for efficient routing.
3. Methodology
3.1. Definitions
@@
We start by briefly introducing the notions of Local Intrinsic Dimensionality (LID) in the words of analysis. A full treatment can be found, for example, in Houle (DBLP:conf/sisap/Houle17).
@@
Let $\mathcal{X}$ be a domain equipped with a distance measure $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}^{+}$. For a reference point $x\in\mathcal{X}$, let $F_{x}(r)=\mathbb{P}(d(x,Y)\leq r)$ denote the cumulative distribution function (CDF) of the distance between $x$ and a random variable $Y$ drawn from the underlying data distribution. The Local Intrinsic Dimensionality (LID) of $x$, denoted as $\text{LID}(x)$, is defined as the intrinsic growth rate of the probability measure within the neighborhood of $x$:
(1) $\text{LID}(x)\triangleq\lim_{r\to 0}\frac{r\cdot F^{\prime}_{x}(r)}{F_{x}(r)}=\lim_{r\to 0}\frac{d\ln F_{x}(r)}{d\ln r},$
provided the limit exists and $F_{x}(r)$ is continuously differentiable for $r>0$.
@@
(3) $F^{\prime}_{x}(r)\approx C\cdot D\cdot r^{D-1}.$
@@
(4) $D\approx\frac{F^{\prime}_{x}(r)}{F_{x}(r)}\cdot r,$
@@
While Eq. 3.1 provides an intuitive closed-form formula for intrinsic dimensionality, in practice we usually do not have access to the true CDF $F_{x}(r)$. Luckily, we can estimate LID from a finite sample of distances from $x$ to its neighbors using Maximum Likelihood Estimation (MLE) as proposed by (NIPS2004_74934548). According to (10.1145/2783258.2783405), LID can be estimated as follows.
@@
Given a reference point $x$ and its $k$-nearest neighbors determined by the distance measure $d$, let $r_{i}=d(x,v_{i})$ denote the distance to the $i$-th nearest neighbor, sorted such that $r_{1}\leq\dots\leq r_{k}$. Following the formulation in (10.1145/2783258.2783405), which adapts the Hill estimator for intrinsic dimensionality, the LID at $x$ is estimated as:
(5) $\widehat{\text{LID}}(x)=-\left(\frac{1}{k}\sum_{i=1}^{k}\ln\frac{r_{i}}{r_{k}}\right)^{-1}.$
3.2. Mapping Function
The primary goal of Manifold-Consistent Graph Indexing is that the graph topology should adapt to the local geometric complexity. In regions where the Local Intrinsic Dimensionality (LID) is low, the data manifold approximates a flat Euclidean subspace. In such isotropic regions, the Euclidean metric is a reliable proxy for geodesic distance, allowing for aggressive edge pruning (larger $\alpha$ in (subramanya2019diskann)) to permit long-range direct connections without risking semantic shortcuts. Conversely, regions with high LID typically exhibit significant curvature, noise, or singularity. In this case, the Euclidean distance often violates the manifold geodesic structure. To preserve topological fidelity, the indexing algorithm must adopt a conservative pruning strategy (smaller $\alpha$ in (subramanya2019diskann)), thereby forcing the search to take smaller, safer steps along the manifold surface.
Let $u\in V$ be a node in the graph, and $\widehat{\text{LID}}(u)$ be its estimated LID. We define the pruning parameter $\alpha(u)$ as:
(6) $\alpha(u)\triangleq\Phi(\widehat{\text{LID}}(u)).$
The function $\Phi:\mathbb{R}^{+}\to[\alpha_{\min},\alpha_{\max}]$ is designed to satisfy the following geometric intuition: in regions with high LID, the graph should enforce a stricter connectivity constraint (smaller $\alpha$) to avoid short-circuiting the manifold; conversely, in low-LID regions, the constraint can be relaxed (larger $\alpha$).
@@ 3.2. Mapping Function
(7) $z(u)=\frac{\widehat{\text{LID}}(u)-\mu_{\widehat{\text{LID}}}}{\sigma_{\widehat{\text{LID}}}}$,
where $\mu_{\widehat{\text{LID}}}$ and $\sigma_{\widehat{\text{LID}}}$ denote the mean and standard deviation of the set of estimated LID values $\{\widehat{\text{LID}}(v)\mid v\in V\}$ computed across the entire graph.
We then formulate $\Phi$ using a logistic function to smoothly map the Z-score to the operational range $\alpha_{\min},\alpha_{\max}]$:
(8) $\Phi(\widehat{\text{LID}}(u))=\alpha_{\min}+\frac{\alpha_{\max}-\alpha_{\min}}{1+\exp(z(u))}.$
We employ the logistic function over a linear mapping to exploit its saturation properties. LID estimates often exhibit heavy-tailed distributions with extreme outliers. A linear mapping would be hypersensitive to these outliers, skewing the $\alpha$ values for the majority of the data. The logistic function acts as a robust soft-thresholding mechanism: it reduces the variance in the high-LID and low-LID tails (saturating towards $\alpha_{\min}$ and $\alpha_{\max}$, respectively) while maintaining sensitivity in the transition region around the population mean. We set $\alpha_{\min}=1.0$ and $\alpha_{\max}=1.5$ following standard practices in graph indexing (subramanya2019diskann), unless otherwise stated. This formulation ensures that nodes with average complexity ($z(u)\approx 0$) are assigned $\alpha\approx 1.25$, while nodes with significantly higher complexity ($z(u)>0$) are penalized with a stricter $\alpha$ approaching the limit of 1.0.
The mapping function $\Phi$ satisfies the following geometric properties that are essential for stable graph construction: Monotonicity and Boundedness.

### v2

@@
Abstract
Graph-based Approximate Nearest Neighbor (ANN) search often suffers from performance degradation in high-dimensional spaces due to the Euclidean-Geodesic mismatch, where greedy routing diverges from the underlying data manifold. To address this challenge, we propose Manifold-Consistent Graph Indexing (MCGI), a geometry-aware and disk-resident indexing method that leverages Local Intrinsic Dimensionality (LID) to dynamically adapt search strategies to the intrinsic geometry of the data. Unlike standard algorithms that treat dimensions uniformly, MCGI modulates its beam search budget based on in situ geometric analysis, eliminating the dependency on static hyperparameters. Theoretical analysis confirms that MCGI provides robust approximation guarantees by preserving manifold-consistent topological connectivity. Extensive evaluations against three industry-standard baselines across five datasets, ranging from million to billion scales, demonstrate the superiority of our approach. Empirically, MCGI achieves 5.8x higher throughput at 95% recall on the high-dimensional GIST1M dataset compared to the state-of-the-art DiskANN. On the billion-scale SIFT1B and T2I-1B datasets, MCGI further validates its scalability by reducing high-recall query latency by 3x, while maintaining performance parity on standard lower-dimensional benchmarks.
1 Introduction
@@
The advent of Large Language Models [4, 37] has fundamentally transformed the landscape of information retrieval and knowledge management. To address the inherent limitations of LLMs, such as hallucinations [19] and knowledge cutoff dates, Retrieval-Augmented Generation (RAG) [25] has emerged as a critical architectural paradigm. RAG relies heavily on the ability to retrieve semantically relevant context from massive corpora in real-time, typically via dense vector representations [22]. This dependency has placed Approximate Nearest Neighbor Search (ANNS) at the core of modern data infrastructure, demanding vector indices that can scale to billion-point datasets [33] while maintaining low latency and high recall under strict production constraints.
@@
State-of-the-art ANNS solutions have noticeably converged on graph-based indices, with DiskANN (Vamana) [34] being a representative example for large-scale, SSD-resident workloads. These algorithms typically employ greedy routing on a proximity graph to navigate from an entry point to the query target. While such methods exhibit exceptional performance on standard benchmarks like SIFT1M [17] (128 dimensions), their efficiency degrades significantly in high-dimensional spaces, such as GIST1M (960 dimensions). This degradation is often attributed to the curse of dimensionality [2], where the distance contrast diminishes, and the Euclidean shortest path on the graph diverges from the geodesic path on the underlying data manifold . We refer to this phenomenon as the Euclidean-Geodesic mismatch. When the routing algorithm ignores the intrinsic geometry of the data, it performs excessive backtracking and disk I/O, rendering the search inefficient, which is particularly problematic in high-recall regimes where production systems usually expect Recall@10 $\geq$ 95%.
@@
Our key insight is that high-dimensional real-world data is rarely uniformly distributed. Instead, it typically adheres to the Manifold Hypothesis [35, 31], residing on lower-dimensional structures embedded within the ambient space. Consequently, the search difficulty is not uniform across the dataset but is modulated by the Local Intrinsic Dimensionality (LID) [24]. In regions where the data manifold is flat (low LID), greedy routing is effective; however, in regions with high curvature or complex topology (high LID), standard greedy strategies fail to identify the correct descent direction. We thus argue that an optimal indexing strategy must be manifold-aware, dynamically allocating computational resources based on the local geometric complexity.
To achieve that, we introduce Manifold-Consistent Graph Indexing (MCGI), a geometry-aware disk-based indexing architecture designed to align Euclidean search with the underlying data manifold. By integrating LID estimation directly into the routing logic, MCGI adapts its traversal strategy to the local topology of the data.
@@
  • We establish a theoretical method linking local intrinsic dimensionality to graph navigability, offering a rigorous justification for adaptive beam search on non-Euclidean manifolds.
@@
  • We develop a lightweight adaptive routing algorithm that dynamically modulates the search budget based on real-time geometric analysis. This design eliminates the dependency on static and manually tuned hyperparameters that often limit the adaptability of existing methods.
  • We conduct comprehensive empirical evaluations against three industry-standard baselines across five datasets ranging from million to billion scales. Results demonstrate that MCGI achieves up to 5.8x higher throughput on high-dimensional vectors (e.g., 960-dimensional GIST1M) and reduces query latency by 3x on the billion-scale SIFT1B compared to the state-of-the-art baselines, confirming its robustness in mitigating both I/O and computational bottlenecks across diverse production environments.
2 Related Work
Vector Indexing Paradigms
While traditional sparse retrieval methods rely on lexical matching [30], memory resident graph indices like HNSW [27] establish state-of-the-art dense retrieval but face billion-scale memory bottlenecks. Disk-based approaches like DiskANN [34] and SPANN [5], along with predecessors like NSG [9], adapt topologies or inverted structures for SSDs. Recent advancements such as CSPG [39] further optimize sparse proximity graphs to reduce unpromising explorations, yet these methods predominantly rely on static routing parameters evaluated on moderate dimensionality. Unlike these static paradigms, MCGI dynamically modulates the search budget using the estimated LID to align graph traversal with the underlying manifold structure.
High-Dimensional Indexing
Early space partitioning trees like the KD-tree [3] and R-tree [12] severely degrade to linear scan performance in dimensions higher than 20 due to the curse of dimensionality. Approximate methods address this through sub-linear hashing like LSH [15, 7], which incurs excessive storage overhead, or subspace quantization like PQ [17] and OPQ [10], and randomized structures like RP-Trees [6], which sacrifice accuracy through compression. Furthermore, recent literature extensively studies the theoretical limits of navigable graphs [8] and the efficiency of index construction [41] in strictly high-dimensional regimes. MCGI circumvents aggressive quantization losses and redundant hashing overheads by preserving the exact connectivity benefits of graph traversal adapted to local geometry.
Intrinsic Dimensionality
Theoretical metrics such as the doubling dimension [11] and expansion dimension [21] provide asymptotic bounds on search complexity, and recent analyses establish worst-case performance limits for popular implementations like DiskANN and HNSW based on intrinsic dimensionality [16]. Maximum likelihood estimators [24] for Local Intrinsic Dimensionality enable robust practical estimations . Existing works typically utilize LID passively for query hardness prediction [13, 38] or detecting adversarial examples [26] without altering the underlying index structure. Diverging from passive descriptive applications, MCGI actively utilizes LID as a prescriptive control signal to dynamically modulate graph traversal parameters for efficient routing.
3 Methodology
3.1 Notations and Definitions
@@
We start by briefly introducing the notions of Local Intrinsic Dimensionality (LID) in the words of analysis. A full treatment can be found, for example, in Houle [14].
@@
Let $\mathcal{X}$ be a domain equipped with a distance measure $d:\mathcal{X}\times\mathcal{X}\to\mathbb{R}^{+}$. For a reference point $x\in\mathcal{X}$, let $F_{x}(r)=\mathbb{P}(d(x,Y)\leq r)$ denote the cumulative distribution function (CDF) of the distance between $x$ and a random variable $Y$ drawn from the underlying data distribution, where $\mathbb{P}$ denotes the probablity function and $r$ denotes the radius. The Local Intrinsic Dimensionality (LID) of $x$, denoted as $\text{LID}(x)$, is defined as the intrinsic growth rate of the probability measure within the neighborhood of $x$:
$\text{LID}(x)\triangleq\lim_{r\to 0}\frac{r\cdot F^{\prime}_{x}(r)}{F_{x}(r)}=\lim_{r\to 0}\frac{d\ln F_{x}(r)}{d\ln r},$ (1)
provided that the limit exists and $F_{x}(r)$ is continuously differentiable for $r>0$.
@@
$F^{\prime}_{x}(r)\approx C\cdot D\cdot r^{D-1}.$ (3)
@@
$D\approx\frac{F^{\prime}_{x}(r)}{F_{x}(r)}\cdot r,$ (4)
@@
While Eq. (1) provides an intuitive closed-form formula for intrinsic dimensionality, in practice we usually do not have access to the true CDF $F_{x}(r)$. Instead, we can estimate LID from a finite sample of distances from $x$ to its neighbors using Maximum Likelihood Estimation (MLE), as proposed by [24]. According to [1], LID can be estimated as follows.
@@
Given a reference point $x$ and its $k$-nearest neighbors determined by the distance measure $d$. Let $r_{i}=d(x,v_{i})$ denote the distance to the $i$-th nearest neighbor, sorted such that $r_{1}\leq\dots\leq r_{k}$. Following the formulation in [1], which adapts the Hill estimator for intrinsic dimensionality, the LID at $x$ can be estimated as:
$\widehat{\text{LID}}(x)=-\left(\frac{1}{k}\sum_{i=1}^{k}\ln\frac{r_{i}}{r_{k}}\right)^{-1}.$ (5)
Remark 3.4 (Intuition behind the LID Estimator).
The rationale behind this estimator stems from modeling the local distance distribution as a power law. Within a sufficiently small neighborhood bounded by $r_{k}$, the cumulative distribution of distances $F_{x}(r)$ grows proportionally to $r^{m}$, where $m$ is the intrinsic dimensionality. By treating the observed normalized distances $r_{i}/r_{k}$ as independent samples drawn from this power law distribution, we can formulate a likelihood function. Solving for the MLE of the exponent $m$ yields exactly the negative inverse of the average log ratios. This elegantly translates the abstract continuous growth rate into a discrete, computationally efficient statistical average.
3.2 Mapping Function from LID to Pruning Parameter
The primary hypothesis of the proposed Manifold-Consistent Graph Indexing (MCGI) is that the graph topology should adapt to the local geometric complexity, rather than being a static vertex-edge structure. While the following formulation may seem technical, MCGI is conceptually straightforward: it uses the estimated LID to determine how aggressively to prune edges during graph construction. That is, in regions where the Local Intrinsic Dimensionality (LID) is low, the data manifold approximates a flat Euclidean subspace. In such isotropic regions, the Euclidean metric is a reliable proxy for geodesic distance, allowing for aggressive edge pruning (larger $\alpha$ in [34]) to permit long-range direct connections without risking semantic shortcuts. Conversely, regions with high LID typically exhibit significant curvature, noise, or singularity. In this case, the Euclidean distance often violates the manifold geodesic structure. To preserve topological fidelity, the indexing algorithm must adopt a conservative pruning strategy (smaller $\alpha$ in [34]), thereby forcing the search to take smaller, safer steps along the manifold surface. The question then becomes how to quantitatively map the estimated LID to the pruning parameter $\alpha$ that controls the edge selection during graph construction. We refer to such a function as the mapping function $\Phi:\mathbb{R}^{+}\to[\alpha_{\min},\alpha_{\max}]$, which translates the geometric complexity of the local neighborhood into a corresponding pruning constraint.
Let $u\in V$ be a node in the graph and $\widehat{\text{LID}}(u)$ be its estimated LID. We define the pruning parameter $\alpha(u)$ as:
$\alpha(u)\triangleq\Phi(\widehat{\text{LID}}(u)),$ (6)
where $\Phi$ will be constructed to satisfy the following geometric requirements: in regions with high LID, the graph should enforce a stricter connectivity constraint (smaller $\alpha$) to avoid short-circuiting the manifold; conversely, in low-LID regions, the constraint can be relaxed (larger $\alpha$).
@@ 3.2 Mapping Function from LID to Pruning Parameter
$z(u)=\frac{\widehat{\text{LID}}(u)-\mu_{\widehat{\text{LID}}}}{\sigma_{\widehat{\text{LID}}}}$, (7)
where $\mu_{\widehat{\text{LID}}}$ and $\sigma_{\widehat{\text{LID}}}$ denote the mean and standard deviation of the set of estimated LID values $\{\widehat{\text{LID}}(v)\mid v\in V\}$ computed across the entire graph. We then formulate $\Phi$ using a logistic function to smoothly map the Z-score to the operational range $\alpha_{\min},\alpha_{\max}]$:
$\Phi(\widehat{\text{LID}}(u))=\alpha_{\min}+\frac{\alpha_{\max}-\alpha_{\min}}{1+\exp(z(u))}.$ (8)
We employ a logistic function rather than a linear mapping to mitigate the impact of extreme outliers inherent in LID estimations. Because LID distributions are typically heavy tailed, a linear scaling would be severely skewed by a few extreme values, compressing the $\alpha$ assignments for the vast majority of nodes into a narrow, indistinguishable band. The logistic function resolves this by acting as a smooth squashing mechanism. Centered on the Z score, it allocates its steepest and most sensitive region to the bulk of the population near the mean. Simultaneously, it safely clamps outlier nodes to the operational boundaries without distorting the global mapping resolution. In experiments (§5), we set $\alpha_{\min}=1.0$ and $\alpha_{\max}=1.5$ following standard practices in graph indexing [34], unless otherwise stated. This formulation ensures that nodes with average complexity ($z(u)\approx 0$) receive an intermediate parameter ($\alpha\approx 1.25$), while those in highly complex regions ($z(u)>0$) are smoothly penalized toward the strict routing limit of 1.0.
The mapping function $\Phi$ satisfies the following geometric properties that are essential for stable graph construction: Monotonicity and Boundedness. We state these properties formally as propositions and provide proofs in the Appendix.